# Multipartite entanglement distribution in a topological photonic network

### v1

@@ III.2 Performance against disorder: concurrence
where each $h_{j,\alpha;j^{\prime},\alpha^{\prime}}^{(\sigma)}$ is an independent random variable sampled once per realization with a normal probability distribution centered at zero and with a standard deviation of $\sigma$, which corresponds to the intensity of disorder. We tune the diagonal and off-diagonal levels of disorder independently to showcase the topological protection of our system towards off-diagonal (OD) symmetry-preserving disorder as well as to check the impact of general (G) symmetry-breaking noise.
There is a second feature of the system with regards to noise protection. In Zurita et al. (2023) it was shown that the use of topological domains drastically accelerates transfer protocols—thus reducing the effects of disorder—and increases connectivity. To clarify this quantitatively, we consider two different kinds of simulations: a single, long SSH chain between the qubits with 14 cavities in a line, and the four-domain lattice shown in Fig. 2 (a), with 15 cavities between the qubits. The transfer time goes from $t_{\mathrm{tr}}=304.0/w$ for one domain (with a preparation time of $t_{\mathrm{prep}}=25/w$) to $t_{\mathrm{tr}}=45.3/w$ for four domains (with $t_{\mathrm{prep}}=20/w$). This speed-up, which is exponential in the length of the chain, is crucial in reducing the effects of diagonal disorder on entanglement production.
We show the results in Fig. 3 (b). In the case of off-diagonal disorder (OD), topological protection helps keep the concurrence greater than $0.7$ even at $\sigma=0.1$ (a disorder of $10\%$ with respect to $w$). The single- and four-domain cases perform very similarly, with the single-domain case being slightly better, given that more edge states participate in the transfer in the four-domain case. However, when general disorder is considered, the four-domain case performs much better than the single-domain one. With the protecting symmetry broken, the deciding factor is now the speed of the protocol, and the presence of domain walls allows it to be over in a much shorter time. For more details on the kind and importance of the observed errors see Appendix D.
@@
Figure 3: (a) Preparation protocols for the $\Phi$-type Bell states. Qudit states are represented as blue (0) or yellow (1). Photonic cavities are represented as thin pink lines, with an additional shading if a photon is present. Superpositions are represented with dotted lines. (b) Average concurrence and its standard deviation over 1000 realizations of the Bell states as a function of off-diagonal (OD) and general (G) disorder. We show the single-domain case (blue) and the four-domain case (red).
To prepare the other kind of Bell state, $|\Psi\rangle=(|00\rangle+\xi|11\rangle)/\sqrt{2}$, the same protocol can be used, with an additional $X$ gate acting on the second qubit at the end. We also explore another possibility, using qutrits, in Appendix E.
IV Multipartite entanglement generation
The full connectivity of the topological photonic lattice opens the door to other protocols that, instead of entangling arbitrary pairs of qubits, entangle triplets or larger subsets of qubits. This approach displays the benefits of this nearly one-dimensional platform when it comes to scaling the number of qubits. To demonstrate this, we design protocols that prepare multipartite entangled states. Unlike in bipartite entanglement, there is no unique notion of maximally entangled multi-qubit states that, once generated with the photonic lattice, allows for the generation of any other state without the use of the photonic lattice. Instead, there are multiple families of states, each of them with a maximal state. Thus, we propose protocols for the generation of these maximal states for different families Gour and Wallach (2013). The simplest of these can be regarded as the W and GHZ states for three qubits Dür et al. (2000).
IV.1 W state preparation
@@
IV.2 GHZ state preparation
To prepare the GHZ state, we need qutrits instead of qubits. We now suppose the gap between states $|1\rangle$ and $|2\rangle$, $\Delta\epsilon_{12}$, can be tuned to the frequency of the cavity. State $|2\rangle$ of each qutrit is used as an auxilliary state and is not present in the final result. The protocol is as follows: (i) Start in state $(|011\rangle+\eta|211\rangle)/\sqrt{2}$ (with $|\eta|=1$), which has no entanglement, and an empty photonic lattice. (ii) Implement a $\pi$-pulse on the left qutrit. (iii) Perform a full transfer along the first domain. (iv) Implement a $\pi$-pulse on the second qutrit. (v) Apply an $X$ gate to that qutrit [^3] We implement operations $X$ and $\mathcal{A}$ using a sine-shaped pulse on the appropriate elements of the qudit hamiltonian: the off-diagonal elements that flip the qutrit states and the diagonal elements that fix the ensuing phase shift. Due to the method used, a global phase of $-i$ is acquired, in addition to the implemented operation (e.g. to do an $X$ operation on the left qubit we actually implement $-i(X_{L}\otimes\mathbbm{1}_{R})$).. (vi) Repeat operations (ii-v) along the second domain: do a $\pi$-pulse on the second qutrit, realize a transfer along the second domain and an $X$ gate on the third qutrit. (vii) Finish the protocol by applying operation $\mathcal{A}=|1\rangle\langle 2|+h.c.$ to the last qutrit [see Fig. 4 (a.2)]. The final state is:
$i|\textrm{GHZ}\rangle=i(|000\rangle+\eta\zeta^{2}|111\rangle)/\sqrt{2}$ (10)
In our case, $\zeta=i$, so we choose $\eta=-1$ to get no relative phase.
IV.3 Performance against disorder: entanglement witnesses
@@
$\rho_{inseparable}=P_{1}\rho_{1}\otimes\rho_{23}+P_{2}\rho_{2}\otimes\rho_{13}+P_{3}\rho_{3}\otimes\rho_{12}$ (11)
where the entanglement, if any, is of bipartite nature, but our information about the system is not enough to know between which qubit pair, nor which is the separable one. The subindices of each density matrix $\rho$ indicate the subsystems they span. Whenever a density matrix cannot be expressed as in Eq. 11, or as incoherent mixtures of these matrices, we claim that its tripartite entanglement is genuine. For the sake of brevity, we use witnesses instead of entanglement measures to make assesments about the presence of entanglement as a function of disorder. Witnesses are functions of the state that, in our case, report the presence of genuine entanglement whenever they are positive. If they are non-positive the state might be entangled, but these quantities do not witness it in those particular settings, hence the name. In other words, witnesses are typically built to be sensitive to a particular state and its surroundings. This is the case for the two witnesses that we consider. The first one is tailored to detect the W state in Eq. 9 and is
@@
where $\rho$ is the final density matrix for one realization of the protocol and $\Pi_{\textrm{W}}$ is the projector on state $|\textrm{W}\rangle_{LCR}$. Moreover, we consider another witness, tailored to the GHZ state with no relative phase in Eq. 10:
@@
where $\Pi_{\textrm{GHZ}}$ is the projector on that state. For the derivation of these witnesses, see Acίn et al. Acín et al. (2001).
Figure 4: (a) Preparation protocols for (a.1) W and (a.2) GHZ states for 3 qudits: $L,C,R$. Qudit states are represented as blue (0), yellow (1) or red (2). Photonic cavities are represented as thin pink lines, with an additional shading if a photon is present. Superpositions are represented with dotted lines. (b) Average values over 1000 realizations of the witnesses in Eq. 12 and Eq. 13 as a function of off-diagonal (blue) and general (red) disorder.
@@
The average values of these witnesses as a function of disorder are presented in Fig. 4 (b). From the data we conclude that topological protection plays a fundamental role in the generation of genuine entanglement, since the witnesses remain positive for higher values of symmetry-preserving disorder (blue) than when general, symmetry-breaking disorder is present (red). Nevertheless, the requirements placed on the system’s disorder are higher with multipartite than with bipartite entanglement, since the former is a more valuable resource than the latter. In the bipartite case, some entanglement is still present when the disorder takes values as high as $0.1w$, whereas here we need a disorder smaller than $0.05w$ to detect multipartite entanglement. However, we must point out that, while concurrence is a measure for two-qubit entanglement, the witnesses used for three-qubit entanglement are not measures. Therefore, it could be the case that multipartite entanglement is still present for disorder higher than $0.05w$ and further developments in the theory of genuine entanglement certification Friis et al. (2018) could detect them. We include the fidelity results for both protocols in Appendix D, and discuss the most common errors observed that harm performance.
@@
On a similar note, we remind the reader that this multipartite entanglement can be produced on demand between any three qubits in the scheme of Fig. 2. Additionally, the topological transfer considered here could entangle more than three qubits in a genuine way, leading to further technological applications while paving the way to become a quantum computing platform. The modifications needed to prepare generalized W and GHZ states for any number of qubits are straightforward, and we discuss them in Appendix F.
V Conclusions
@@
In this work, we propose a realistic experimental setting in which fast multidomain topological protocols can be used to quickly and reliably generate multipartite entanglement. We consider an easily scalable photonic lattice in silicon, with coupled qudits implemented natively with impurities, but our protocols could also be considered in other systems like superconducting or spin qubits, both of which are being investigated as platforms for photon-mediated entanglement generation Yang et al. (2016); Chan et al. (2023).
@@ Appendix A Topological protection
In order to prove the topological protection of the model analytically, we first look at the case where all $v$ terms are nonzero but smaller than $w$. As noted recently in Dias and Marques (2022), site indexing plays a vital role in the topological description of 1D models. By choosing the appropriate indexing [which we also use in Eq. (2)], we can conceptualize the SSSH lattice as a series of topological domains with alternating winding numbers of $\nu=\pm 1$. We show this in Fig. 7 (a) for a 5-domain model. The odd-numbered domains are regular SSH chains with a winding number of $1$, while the even domains have a winding number of $-1$. Given that each site connects to the contiguous unit cell that is further from it, instead of closer like the SSH chain, the bulk Hamiltonian $\mathcal{H}(k)$ is identical to the SSH chain but with $k\to-k$. That reverses the path traced by $\mathcal{H}(k)$ in the space of possible Hamiltonians, changing the sign of its winding number, as shown in Fig. 7 (b). The model we use here is completely equivalent to the one in the main text, as shown in Fig. 7 (c). The latter does not use long-range hopping terms.
@@
During a multidomain transfer, the intermediate qubit cavities are left unconnected, as shown in Fig. 2 (d). This case can be seen as a multidomain SSH chain, the topology of which has been discussed in the literature at length Su et al. (1980); Munoz et al. (2018); Qi et al. (2021); Zurita et al. (2023). The topological protection is also guaranteed under this point of view.
@@
This small energy, however, is the cause of the exponentially long durations achieved with single-domain chains Zurita et al. (2023). To fight this, we increase the number of domains, making them closer to each other. As a consequence, their hybridization energies are dramatically larger, as shown in Fig. 5 (c) for the four-domain, 15-cavity SSH model corresponding to the connected sites in Fig. 2 (d). They are still clearly separated from the bulk, but their energy fluctuations are larger than in the single-domain case.
@@
We show the results in Fig. 6 for the single-domain, 14-site model (a,b) and the four-domain, 15-site one (c,d), for both kinds of disorder and $10^{4}$ realizations. Each of the realizations has a fixed disorder profile, scaled to match $\sigma$. As expected, all edge states show a smaller $\sigma_{\epsilon}$ than the bulk states for off-diagonal disorder, although the single-domain case fluctuates much less, as expected. In the general disorder case, however, the single-domain model end modes inmediately lose all protection, while the four-domain model states continue being the more constant ones. This shows that, for a fixed length, the topological protection of a multidomain model is weaker than the single-domain case but, in the more realistic case of general disorder, can resist for larger disorder levels.
The other hallmark of topological protection, bulk-boundary decoupling, is studied in depth in Appendix D, where we break down the results of all disorder simulations.

### v2

@@ 3.2 Performance against disorder: concurrence
where each $h_{j,\alpha;j^{\prime},\alpha^{\prime}}^{(\sigma)}$ is an independent random variable sampled once per realization with a normal probability distribution centered at zero and with a standard deviation of $\sigma$, which corresponds to the intensity of disorder. We tune the diagonal and off-diagonal levels of disorder independently to showcase the topological protection of our system towards off-diagonal (OD) symmetry-preserving disorder, as well as to check the impact of general (G) symmetry-breaking noise.
There is a second feature of the system with regard to noise protection. In [19] it was shown that the use of topological domains drastically accelerates transfer protocols—thus reducing the effects of disorder—and increases connectivity. To clarify this quantitatively, we consider two different kinds of simulations: a single, long SSH chain between the qubits with 14 cavities in a line, and the four-domain lattice shown in Fig. 2 (a), with 15 cavities between the qubits. The transfer time goes from $t_{\mathrm{tr}}=304.0/w$ for one domain (with a preparation time of $t_{\mathrm{prep}}=25/w$) to $t_{\mathrm{tr}}=45.3/w$ for four domains (with $t_{\mathrm{prep}}=20/w$). This speed-up, which is exponential in the length of the chain, is crucial in reducing the effects of diagonal disorder on entanglement production.
We show the results in Fig. 3 (b). In the case of off-diagonal disorder (OD), topological protection helps keep the concurrence greater than $0.7$ even at $\sigma=0.1$ (a disorder of $10\%$ with respect to $w$). The single- and four-domain cases perform very similarly, with the single-domain case being slightly better, given that more edge states participate in the transfer in the four-domain case. However, when general disorder is considered, the four-domain case performs much better than the single-domain one. With the protecting symmetry broken, the deciding factor is now the speed of the protocol, and the presence of domain walls allows it to complete in a much shorter time. For more details on the kind and importance of the observed errors see Appendix E.
@@
Figure 3: (a) Preparation protocols for the $\Phi$-type Bell states. Qudit states are represented as blue (0) or yellow (1). Photonic cavities are represented as thin pink lines, with an additional shading if a photon is present. Superpositions are represented with dotted lines. (b) Average concurrence and its standard deviation over 1000 realizations of the Bell states as a function of off-diagonal (OD) and general (G) disorder. We show the single-domain case ($N_{cav}=14$, blue) and the four-domain case ($N_{cav}=15$, red).
To prepare the other kind of Bell state, $|\Psi\rangle=(|00\rangle+\xi|11\rangle)/\sqrt{2}$, the same protocol can be used, with an additional $X$ gate acting on the second qubit at the end. We also explore another possibility, using qutrits, in Appendix F.
4 Multipartite entanglement generation
The full connectivity of the topological photonic lattice opens the door to other protocols that, instead of entangling arbitrary pairs of qubits, entangle triplets or larger subsets of qubits. This approach displays the benefits of this nearly one-dimensional platform when it comes to scaling the number of qubits. To demonstrate this, we design protocols that prepare multipartite entangled states. Unlike in bipartite entanglement, there is no unique notion of maximally entangled multi-qubit states that, once generated with the photonic lattice, allows for the generation of any other state without the use of the photonic lattice. Instead, there are multiple families of states, each of them with a maximal state. Thus, we propose protocols for the generation of these maximal states for different families [59]. The simplest of these can be regarded as the W and GHZ states for three qubits [60].
4.1 W state preparation
We propose a protocol for the preparation of the W state that is similar to the previous one preparing the $\Phi$ state in Eq. (6), but with extended notation regarding the third qubit, which we label $C$ after its centered position somewhere between the leftmost and rightmost qubits. The setup is shown in Fig. 4 (a). With this extension in notation, the protocol is as follows.
@@ 4.1 W state preparation
Figure 4: (a) Preparation protocols for the (a.1) W and (a.2) GHZ states for 3 qudits: $L,C,R$. Qudit states are represented as blue (0), yellow (1) or red (2). Photonic cavities are represented as thin pink lines, with an additional shading if a photon is present. Superpositions are represented with dotted lines. (b) Average values over 1000 realizations of the W witness in Eq. (12) (b.1) and the GHZ witness in Eq. (13) (b.2) as a function of off-diagonal (orange) and general (purple) disorder. The lattice has $N=2$ domains and $N_{cav}=10$ cavities.
4.2 GHZ state preparation
To prepare the GHZ state, we need qutrits instead of qubits. We now suppose the gap between states $|1\rangle$ and $|2\rangle$, $\Delta\epsilon_{12}$, can be tuned to the frequency of the cavity. State $|2\rangle$ of each qutrit is used as an auxiliary state and is not present in the final result. The protocol is as follows.
(i) Start in state $(|011\rangle+\eta|211\rangle)/\sqrt{2}$ (with $|\eta|=1$), which has no entanglement, and an empty photonic lattice. (ii) Implement a $\pi$-pulse on the left qutrit. (iii) Perform a full transfer along the first domain. (iv) Implement a $\pi$-pulse on the second qutrit. (v) Apply an $X$ gate to that qutrit [^3] We implement operations $X$ and $\mathcal{A}$ using a sine-shaped pulse on the appropriate elements of the qudit Hamiltonian: the off-diagonal elements that flip the qutrit states and the diagonal elements that fix the ensuing phase shift. Due to the method used, a global phase of $-i$ is acquired, in addition to the implemented operation (e.g. to do an $X$ operation on the left qubit we actually implement $-i(X_{L}\otimes\mathbbm{1}_{R})$).. (vi) Repeat operations (ii-v) along the second domain: do a $\pi$-pulse on the second qutrit, realize a transfer along the second domain, and a $\pi$-pulse and $X$ gate on the third qutrit. (vii) Finish the protocol by applying operation $\mathcal{A}=|1\rangle\langle 2|+h.c.$ to the last qutrit [see Fig. 4 (a.2)]. The final state is:
$i|\textrm{GHZ}\rangle=i(|000\rangle+\eta\zeta^{2}|111\rangle)/\sqrt{2}.$ (10)
In our case, $\zeta=i$, so we choose $\eta=-1$ to get no relative phase. We include the intermediate states for this and all protocols in Appendix D.
4.3 Performance against disorder: entanglement witnesses
@@
$\rho_{\textrm{insep}}=P_{1}\rho_{1}\otimes\rho_{23}+P_{2}\rho_{2}\otimes\rho_{13}+P_{3}\rho_{3}\otimes\rho_{12}$ (11)
where the entanglement, if any, is of bipartite nature, but our information about the system is not enough to know between which qubit pair, nor which is the separable one. The subindices of each density matrix $\rho$ indicate the subsystems they span. Whenever a density matrix cannot be expressed as in Eq. (11), or as incoherent mixtures of these matrices, we claim that its tripartite entanglement is genuine. For the sake of brevity, we use witnesses instead of entanglement measures to make assessments about the presence of entanglement as a function of disorder. Witnesses are functions of the state that, in our case, report the presence of genuine entanglement whenever they are positive. If they are non-positive the state might be entangled, but these quantities do not witness it in those particular settings, hence the name. In other words, witnesses are typically built to be sensitive to a particular state and its surroundings. This is the case for the two witnesses that we consider. The first one is tailored to detect the W state in Eq. (9) and is
@@
where $\rho$ is the final density matrix for one realization of the protocol, and $\Pi_{\textrm{W}}$ is the projector on state $|\textrm{W}\rangle_{LCR}$. Moreover, we consider another witness, tailored to the GHZ state with no relative phase in Eq. (10):
@@
where $\Pi_{\textrm{GHZ}}$ is the projector on that state. For the derivation of these witnesses, see [61].
@@
The average values of these witnesses as a function of disorder are presented in Fig. 4 (b), for lattice with 10 cavities, 2 domains ($\ell=4$) and 3 qudits. From the data we conclude that topological protection plays a fundamental role in the generation of genuine entanglement, since the witnesses remain positive for higher values of symmetry-preserving disorder (purple) than when general, symmetry-breaking disorder is present (orange). Nevertheless, the requirements placed on the system’s disorder are higher with multipartite than with bipartite entanglement, since the former is a more valuable resource than the latter. In the bipartite case, some entanglement is still present when the disorder takes values as high as $0.1w$, whereas here we need a disorder smaller than $0.05w$ to detect multipartite entanglement. However, we must point out that, while concurrence is a measure for two-qubit entanglement, the witnesses used for three-qubit entanglement are not measures. Therefore, it could be the case that multipartite entanglement is still present for disorder higher than $0.05w$ and further developments in the theory of genuine entanglement certification [62] could detect them. We include the fidelity results for both protocols in Appendix E, and discuss the most common errors observed that harm performance.
@@
On a similar note, we remind the reader that this multipartite entanglement can be produced on demand between any three qubits in the scheme of Fig. 2. Additionally, the topological transfer considered here could entangle more than three qubits in a genuine way, leading to further technological applications while paving the way to become a quantum computing platform. The modifications needed to prepare generalized W and GHZ states for any number of qubits are straightforward, and we discuss them in Appendix G.
5 Probable sources of loss and decoherence
Considering the experimental platform detailed in Section 2.3, we now explore the main sources of loss and decoherence that might appear in the qudits and photonic lattice, and assess their effects on the protocols.
5.1 Spontaneous qubit deexcitation
The main source of decoherence affecting the color center qudits is spontaneous deexcitation, in which a qudit decays to the ground state without emitting a photon in the cavity. As discussed in similar setups [63], this causes the amplitude of the excited qudit states to decay in time. The fidelity of the final state then shows an exponential decay:
$f=f_{0}e^{-\gamma t_{\textrm{tr}}},$ (14)
where $f_{0}$ is the fidelity with no decoherence, $t_{\textrm{tr}}$ is the duration of the protocol and $\gamma$ is the spontaneous qubit deexcitation rate.
[…]
5.2 Photonic scattering losses
The sources of photonic loss in the cavity lattice can be multiple. Radiative losses, in which the photon in a cavity is lost because of radiation, are experimentally found to be negligible in air-hole cavity lattices in semiconductors [65]. Scattering losses, caused by imperfections in the air-hole lattice, are much more important. These imperfections can be perturbations on the air-hole radii, in their positions, or in the angle their walls form with the vertical.
As shown in the literature [65, 66], small deviations in this angle, caused by an imperfect etching process, can substantially alter the performance of the photonic lattice. Any amount of asymmetry in the vertical axis, such as the one caused by non-vertical air-holes, will couple the resonating modes of the cavities with the bulk photonic modes of the surrounding air-hole lattice. This will cause unwanted power losses into those modes, which will end up dissipating.
However, this loss will only be significant if any of the frequencies involved in the transfer is close to a bulk mode frequency of the air-hole lattice [66]. A partial solution would be to keep the transfer frequencies—which are tuned by the control parameter $v$—lower than the bulk mode band, although sacrificing some speed in the process. In any case, the frequency of the cavities will depend strongly on the color centers used and the photonic crystal architecture, and could be more challenging to adjust.
5.3 Off-diagonal disorder between cavities
Finally, there is another source of decoherence, which is specific to the platform we propose. As mentioned in [43, 46], controlling the exact placement and orientation of the color centers can pose a challenge, although advances are being made with controllable growth techniques [43, 44, 46].
A possible solution would be to create the cavity lattice around the impurities after they are placed [46]. The uncertainty in their location would then induce perturbations in the position of each cavity and the distance between them, causing off-diagonal disorder. This disorder in the photonic lattice translates as a source of decoherence in the qubit subspace, as explored in Appendix E.
However, the topological robustness against off-diagonal disorder in our protocols would greatly dampen its effects, as we show in Sections 3 and 4. Thus, these protocols might be especially helpful in the implementation of the model.
6 Conclusions
@@
In this work, we propose a realistic experimental setting in which fast multidomain topological protocols can be used to quickly and reliably generate multipartite entanglement. We consider an easily scalable photonic lattice in silicon, with coupled qudits implemented natively with impurities, but our protocols could also be considered in other systems like superconducting or spin qubits, both of which are being investigated as platforms for photon-mediated entanglement generation [67, 68].
@@ Appendix A Topological protection
In order to prove the topological protection of the model analytically, we first look at the case where all $v$ terms are nonzero but smaller than $w$. As noted recently in [69], site indexing plays a vital role in the topological description of 1D models. By choosing the appropriate indexing [which we also use in Eq. (2)], we can conceptualize the SSSH lattice as a series of topological domains with alternating winding numbers of $\nu=\pm 1$. We show this in Fig. 7 (a) for a 5-domain model. The odd-numbered domains are regular SSH chains with a winding number of $1$, while the even domains have a winding number of $-1$. Given that each site connects to the contiguous unit cell that is further from it, instead of closer like the SSH chain, the bulk Hamiltonian $\mathcal{H}(k)$ is identical to the SSH chain but with $k\to-k$. That reverses the path traced by $\mathcal{H}(k)$ in the space of possible Hamiltonians, changing the sign of its winding number, as shown in Fig. 7 (b). The model we use here is completely equivalent to the one in the main text, as shown in Fig. 7 (c). The latter does not use long-range hopping terms.
@@
During a multidomain transfer, the intermediate qubit cavities are left unconnected, as shown in Fig. 2 (d). This case can be seen as a multidomain SSH chain, the topology of which has been discussed in the literature at length [70, 71, 72, 19]. The topological protection is also guaranteed under this point of view.
@@
This small energy, however, is the cause of the exponentially long durations achieved with single-domain chains [19]. To fight this, we increase the number of domains, making them closer to each other. As a consequence, their hybridization energies are dramatically larger, as shown in Fig. 5 (c) for the four-domain, 15-cavity SSH model corresponding to the connected sites in Fig. 2 (d). They are still clearly separated from the bulk, but their energy fluctuations are larger than in the single-domain case.
@@
We show the results in Fig. 6 for the single-domain, 14-site model (a,b) and the four-domain, 15-site one (c,d), for both kinds of disorder and $10^{4}$ realizations. Each of the realizations has a fixed disorder profile, scaled to match $\sigma$. As expected, all edge states show a smaller $\sigma_{\epsilon}$ than the bulk states for off-diagonal disorder, although the single-domain case fluctuates much less, as expected. In the general disorder case, however, the single-domain model end modes immediately lose all protection, while the four-domain model states continue being the more constant ones. This shows that, for a fixed length, the topological protection of a multidomain model is weaker than the single-domain case but, in the more realistic case of general disorder, can resist for larger disorder levels.
The other hallmark of topological protection, bulk-boundary decoupling, is studied in depth in Appendix E, where we break down the results of all disorder simulations.